# Fast Ewald Summation using Prolate Spheroidal Wave Functions

### v1

@@ 5.1 Ewald-split truncation errors
$\frac{1}{V}\,\|\phi^{\mathrm{far}}-\phi^{\mathrm{far}}_{t}\|^{2}_{L^{2}(\Omega)}\;\leq\;\frac{4\,C_{\rho}C_{*}\,r_{c}\,\|\bm{\rho}\|_{2}^{2}}{V\,c_{s}^{2}\,\psi_{0}^{c_{s}}(0)^{2}}\left(\frac{2\pi}{\lambda_{0}^{2}}-c_{s}\right),$ (80)
where $V=L^{3}$ and $0<C_{\rho}\leq n$ is the constant from Lemma 3.
Proof.
By Lemma 4,
$\frac{1}{V}\,\|\phi^{\mathrm{far}}-\phi^{\mathrm{far}}_{t}\|^{2}_{L^{2}(\Omega)}\leq\frac{8\,C_{\rho}}{V}\,\|\bm{\rho}\|_{2}^{2}\int_{\omega_{*}}^{\infty}\frac{|\hat{\gamma}_{r_{c}}^{c_{s}}(\omega)|^{2}}{\omega^{2}}\,d\omega.$ (81)
Splitting the integral at $\omega_{\max}$ and using that $\omega\mapsto\omega^{-2}$ is decreasing,
$\int_{\omega_{*}}^{\infty}\frac{|\hat{\gamma}_{r_{c}}^{c_{s}}(\omega)|^{2}}{\omega^{2}}\,d\omega\leq\frac{1}{\omega_{*}^{2}}\int_{\omega_{*}}^{\omega_{\max}}|\hat{\gamma}_{r_{c}}^{c_{s}}(\omega)|^{2}\,d\omega+\frac{1}{\omega_{\max}^{2}}\int_{\omega_{\max}}^{\infty}|\hat{\gamma}_{r_{c}}^{c_{s}}(\omega)|^{2}\,d\omega.$ (82)
Since $\int_{\omega_{*}}^{\omega_{\max}}\leq\int_{0}^{\omega_{\max}}$, we relate the in-band and tail energies using the PSWF scaling
$\hat{\gamma}_{r_{c}}^{c_{s}}(\omega)=\frac{1}{\lambda_{0}\,\psi_{0}^{c_{s}}(0)}\,\hat{\psi}_{0}^{c_{s}}(u)\big|_{u=r_{c}\omega}.$ (83)
Hence,
$\frac{\displaystyle\int_{0}^{\omega_{\max}}|\hat{\gamma}_{r_{c}}^{c_{s}}(\omega)|^{2}\,d\omega}{\displaystyle\int_{\omega_{\max}}^{\infty}|\hat{\gamma}_{r_{c}}^{c_{s}}(\omega)|^{2}\,d\omega}=\frac{\displaystyle\int_{0}^{c_{s}}|\hat{\psi}_{0}^{c_{s}}(u)|^{2}\,du}{\displaystyle\int_{c_{s}}^{\infty}|\hat{\psi}_{0}^{c_{s}}(u)|^{2}\,du}=\frac{c_{s}\,\lambda_{0}^{2}}{\,2\pi-c_{s}\,\lambda_{0}^{2}\,}$, (84)
where we used Theorem 1 and evenness. Therefore,
$\int_{\omega_{*}}^{\infty}\frac{|\hat{\gamma}_{r_{c}}^{c_{s}}(\omega)|^{2}}{\omega^{2}}\,d\omega\leq\left(\left(\frac{\omega_{\max}}{\omega_{*}}\right)^{\!2}\frac{c_{s}\,\lambda_{0}^{2}}{\,2\pi-c_{s}\,\lambda_{0}^{2}\,}+1\right)\frac{1}{\omega_{\max}^{2}}\int_{\omega_{\max}}^{\infty}|\hat{\gamma}_{r_{c}}^{c_{s}}(\omega)|^{2}\,d\omega.$ (85)
With the change of variables $u=r_{c}\omega$ and $\omega_{\max}=c_{s}/r_{c}$, and substituting $\hat{\gamma}^{c_{s}}_{r_{c}}(\omega)=\psi_{0}^{c_{s}}(r_{c}\omega/c_{s})/\psi_{0}^{c_{s}}(0)$, the tail integral becomes
$\int_{\omega_{\max}}^{\infty}|\hat{\gamma}_{r_{c}}^{c_{s}}(\omega)|^{2}\,d\omega$ $=\frac{1}{\lambda_{0}^{2}\,\psi_{0}^{c_{s}}(0)^{2}\,r_{c}}\int_{c_{s}}^{\infty}|\hat{\psi}_{0}^{c_{s}}(u)|^{2}\,du$
$=\frac{1}{\lambda_{0}^{2}\,\psi_{0}^{c_{s}}(0)^{2}\,r_{c}}\,\frac{1}{2}\bigl(2\pi-c_{s}\,\lambda_{0}^{2}\bigr),$ (86)
again using Theorem 1 and evenness.
Combining these identities yields
$\int_{\omega_{*}}^{\infty}\frac{|\hat{\gamma}_{r_{c}}^{c_{s}}(\omega)|^{2}}{\omega^{2}}\,d\omega\leq\frac{C_{*}\,r_{c}}{2\,c_{s}^{2}\,\psi_{0}^{c_{s}}(0)^{2}}\left(\frac{2\pi}{\lambda_{0}^{2}}-c_{s}\right).$ (87)
Multiplying by $\tfrac{8\,C_{\rho}}{V}\,\|\bm{\rho}\|_{2}^{2}$ gives the claimed bound. ∎
@@
The leading PSWF eigenvalue depends on the bandlimit, $\lambda_{0}=\lambda_{0}(c_{s})$. It is known (see, e.g., [24, p. 42]) that, as $c_{s}\to\infty$,
$\frac{2\pi}{\lambda_{0}^{2}}\sim c_{s}.$ (88)
Hence
$\frac{2\pi}{c_{s}\lambda_{0}^{2}}-1=\frac{1}{c_{s}}\Big(\frac{2\pi}{\lambda_{0}^{2}}\Big)-1\;\to\;0\quad\text{as }c_{s}\to\infty.$ (89)
This convergence is exponentially fast (see Appendix A and [29, 9]). However, numerically this expression is the difference of two nearly equal terms, which leads to catastrophic cancellation in double precision. As a result, $2\pi/(c_{s}\lambda_{0}^{2})-1$ cannot be evaluated reliably for $c_{s}\gtrsim 17$. Using the known asymptotic behavior of $\lambda_{0}$, this quantity can instead be evaluated via a stable and accurate approximation (see Figure 9).
@@ 5.1.1 A closed form approximation for the split truncation error
Although the bound in Proposition 1 is rigorous, it cannot be evaluated reliably for large $c_{s}$ due to subtractive cancellation (see Remark 11). Moreover, its asymptotic decay is not immediately useful for parameter selection. For practical purposes, we therefore introduce an exponential approximation, fitted to numerical data, that captures the decay of the error with high accuracy over the parameter ranges of interest.
Assuming the particle positions are well distributed (i.e., without significant clustering), we may set $C_{\rho}\approx 1$ as in Remark 10. Furthermore, the prefactor $C_{*}$ in Proposition 1 accounts for the contribution from the interval $\omega_{*},\omega_{\max}]$. In practical settings, where the truncation parameter $K_{\text{max}}$ is sufficiently large (so that $\omega_{*}\approx\omega_{\max}$), this factor satisfies $C_{*}\approx 1$ and can therefore be neglected. Remark 11 together with the curve-fit in Table 5 (Appendix A) shows that, for $c_{s}\gtrsim 10$, the expression
$\frac{2}{c_{s}\,\psi_{0}^{c_{s}}(0)}\sqrt{\frac{2\pi}{\lambda_{0}^{2}}-c_{s}}$ (90)
is well approximated (up to a maximal relative error of $\approx 0.2\%$) by an exponential (see Equation (143) in Appendix A). This motivates the following closed form model
$\frac{1}{\sqrt{V}}\,\|\phi^{\mathrm{far}}-\phi^{\mathrm{far}}_{t}\|_{L^{2}(\Omega)}\;\approx\;\frac{1}{\sqrt{V}}\|\bm{\rho}\|_{2}\,\sqrt{r_{c}}\;A_{s}\;c_{s}^{-1/2}\,e^{-c_{s}},\qquad A_{s}=6.91.$ (91)
Finally, we stress that the assumption $C_{\rho}\approx 1$ is essential: for strongly clustered particle configurations, the error may increase by up to a factor of $\sqrt{n}$.
@@ 5.2 Aliasing errors due to window function approximations
The sources of error arising from the NUFFT components are analogous to those in fast Ewald summation. In practice, all window functions are truncated. As discussed in Section 4.3.1 for the Gaussian window, the use of the untruncated Fourier transform introduces additional errors in the spreading and interpolation steps. These are commonly referred to as truncation errors. A second source of error arises from aliasing, which is caused by discrete sampling of the window function. Sampling induces a periodization in Fourier space via the Poisson summation formula, leading to overlapping frequency contributions [25, p. 124].
@@
Since $c_{\boldsymbol{k}}(\tilde{\varphi}_{t})$ is even and monotonically decreasing in $|\boldsymbol{k}|$ for $|\boldsymbol{k}|>0$, there exist indices $\boldsymbol{k}\in\mathcal{I}_{\boldsymbol{m}}$ and $\boldsymbol{r}\in\mathbb{Z}^{3}\setminus\{\boldsymbol{0}\}$ such that $|\boldsymbol{k}+m\boldsymbol{r}|\leq|\boldsymbol{k}|$, implying
$\frac{c^{2}_{\boldsymbol{k}+m\boldsymbol{r}}(\tilde{\varphi}_{t})}{c^{2}_{\boldsymbol{k}}(\tilde{\varphi}_{t})}\geq 1.$ (100)
Hence, in the absence of oversampling, aliasing errors are not uniformly suppressed by the window function alone.
Nevertheless, the diagonal scaling in fast Ewald methods—contained in the $|\hat{f}_{\bm{k}}|^{2}$ factor—damps this effect.
@@ 5.2.1 A closed-form approximation for the aliasing error
In (103), the aliasing sum is dominated by the first nonzero terms with $|\bm{r}|=1$, since the PSWF window decays rapidly with $|\bm{r}|$. These six contributions—corresponding to one shift in each coordinate direction in three dimensions—provide the leading-order contribution to the aliasing error. We therefore approximate (103) by retaining only these terms, yielding
@@
Since $|\bm{k}|\geq|k_{i}|$ and $\psi_{0}^{c}$ is even and nonincreasing on $0,\infty)$, we have
@@
where the approximation uses $c_{s}\approx c_{w}$ with $c_{s}<c_{w}$ (see Section 7).
Next, for $k_{i}\in\mathcal{I}_{m}$ and $r_{i}\in\{\pm 1\}$,
$\left|\tfrac{2}{m}(k_{i}+mr_{i})\right|\in[1,3],\qquad\left|\tfrac{2}{m}k_{i}\right|\in(0,1],$ (111)
and by evenness and monotonicity of $\psi_{0}^{c_{w}}$,
@@
Collecting terms, we obtain
$\sum_{\bm{k}\in\mathcal{I}_{\bm{m}}\setminus\{\bm{0}\}}\frac{\displaystyle\sum_{|\bm{r}|=1}\prod_{i=1}^{3}\psi_{0}^{c_{w}}\!\Big(\tfrac{2}{m}(k_{i}+mr_{i})\Big)^{2}}{|\bm{k}|^{4}\displaystyle\prod_{i=1}^{3}\psi_{0}^{c_{w}}\!\left(\tfrac{2}{m}k_{i}\right)^{2}}$ $\approx 6\,\psi_{0}^{c_{w}}(1)^{2}\sum_{k=1}^{K_{\max}}\frac{1}{k^{4}}\;<\;\psi_{0}^{c_{w}}(1)^{2}\,\frac{\pi^{4}}{15},$ (113)
@@
Substituting (113) into (108) yields the dimensionless prefactor $C_{\rho}\pi^{2}/15\approx 0.66\,C_{\rho}$, which is of order unity and independent of $m$, $L$, and $c_{w}$. We thus arrive at the approximation
$\frac{1}{\sqrt{V}}\|\phi_{t}^{\mathrm{far}}-\phi_{t,h}^{\mathrm{far}}\|_{L^{2}(\Omega)}\;\approx\;\frac{1}{\sqrt{V}}\frac{\pi}{\sqrt{15}}\,\sqrt{C_{\rho}\,L}\,\|\bm{\rho}\|_{2}\,\frac{\psi_{0}^{c_{w}}(1)}{\psi_{0}^{c_{w}}(0)}.$ (114)
Assuming well-distributed particles so that $C_{\rho}\approx 1$, and approximating $\psi_{0}^{c_{w}}(1)/\psi_{0}^{c_{w}}(0)$ accurately (up to a maximal relative error of $\approx 1.5$ %) by curve-fitting (see Table 5 in Appendix A), yields the final closed form model for the aliasing error
$\frac{1}{\sqrt{V}}\|\phi_{t}^{\mathrm{far}}-\phi_{t,h}^{\mathrm{far}}\|_{L^{2}(\Omega)}\;\approx\;\frac{1}{\sqrt{V}}A_{w}\,\sqrt{L}\,\|\bm{\rho}\|_{2}\,c_{w}^{1/2}\,e^{-c_{w}},\qquad A_{w}\approx 2.78.$ (115)
Numerical experiments show that both the split-error model (91) and the aliasing-error model (115) closely match the observed absolute RMS errors produced by Algorithm 2; see Figure 4.
@@ 6 Parameter selection
This section presents practical guidelines for selecting the PSWF mollifier and window parameters to achieve a prescribed tolerance $\varepsilon$. The results rely on the error models of Section 5. Unless otherwise noted, we assume neutral, well-distributed particle configurations so that $C_{\rho}\approx 1$ (cf. Remark 10), no oversampling, and parameters $c_{s},c_{w}$ within the fitted ranges of Table 5.
@@ 6.2 PSWF split parameters
Two parameters control the Ewald split: the shape parameter $c_{s}$ and the cutoff $r_{c}$. The cutoff balances the real- and Fourier-space costs, while $c_{s}$ determines the number of Fourier modes. They are related by (see (55))

### v2

@@ 5.1 Kernel split truncation errors
$=\frac{8Vc_{s}^{2}}{r_{c}^{2}}\left(\frac{\psi_{0}^{c_{s}}(1)}{\psi_{0}^{c_{s}}(0)}\right)^{2}\int_{\omega_{\max}-\tfrac{\sqrt{3}\pi}{L}}^{\infty}\frac{1}{\omega^{4}}\,\mathrm{d}\omega$
$\leq\frac{8Vc_{s}^{2}}{3r_{c}^{2}}\left(\frac{\psi_{0}^{c_{s}}(1)}{\psi_{0}^{c_{s}}(0)}\right)^{2}\frac{1}{(\omega_{\max}-\tfrac{\sqrt{3}\pi}{L})^{3}}.$ (83)
In the equality step we passed to spherical coordinates. Since $\omega_{\max}=c_{s}/r_{c}$, a Taylor expansion yields
$\frac{1}{(\omega_{\max}-\tfrac{\sqrt{3}\pi}{L})^{3}}=\frac{r_{c}^{3}}{c_{s}^{3}}\left(1+\mathcal{O}\!\left(\frac{r_{c}}{c_{s}}\right)\right).$ (84)
Substituting this estimate into (83) gives
$\sum_{|\bm{k}|>K_{\text{max}}}|\hat{M}(\bm{\omega}(\bm{k}))|^{2}\leq\frac{8Vc_{s}^{2}}{3r_{c}^{2}}\left(\frac{\psi_{0}^{c_{s}}(1)}{\psi_{0}^{c_{s}}(0)}\right)^{2}\frac{r_{c}^{3}}{c_{s}^{3}}\left(1+\mathcal{O}\!\left(\frac{r_{c}}{c_{s}}\right)\right).$ (85)
Finally, inserting this bound into (79) completes the proof. ∎
@@
A more refined heuristic than the bound (77) is obtained by assuming that
$|\psi_{0}^{c_{s}}(s)|\approx\frac{2\psi_{0}^{c_{s}}(1)}{\lambda_{0}}\frac{|\sin(c_{s}s)|}{c_{s}s},\qquad s>1,$ (86)
as suggested by Lemma 5. Using $|\sin(c_{s}s)|\leq 1$ and the lower bound $\lambda_{0}>\sqrt{2\pi/c_{s}}$ yields
$|\psi_{0}^{c_{s}}(s)|\lesssim\sqrt{\frac{2}{\pi c_{s}}}\frac{\psi_{0}^{c_{s}}(1)}{s},\qquad s>1.$ (87)
Replacing $|\psi_{0}^{c_{s}}(s)|\leq\psi_{0}^{c_{s}}(1)/s$ by this heuristic estimate in the proof of Proposition 1 leads to the approximation
$\frac{1}{V}\|\phi^{\mathrm{far}}-\phi_{t}^{\mathrm{far}}\|_{L^{2}(\Omega)}^{2}\approx\frac{16\,C_{\rho}\|\bm{\rho}\|_{2}^{2}\,r_{c}}{3\pi V\,c_{s}^{2}}\left(\frac{\psi_{0}^{c_{s}}(1)}{\psi_{0}^{c_{s}}(0)}\right)^{2}.$ (88)
Compared with (77), the heuristic estimate (88) gains an additional factor of $c_{s}^{-1}$.
@@ 5.1.1 A closed form approximation for the split truncation error
For practical parameter selection, we use the heuristic model (88). We approximate the ratio $\psi_{0}^{c_{s}}(1)/\psi_{0}^{c_{s}}(0)\approx A_{2}\,c_{s}^{1/2}e^{-c_{s}}$, where $A_{2}\approx 3.42$, obtained in (145) (see Appendix A). This yields
$\frac{1}{\sqrt{V}}\|\phi^{\mathrm{far}}-\phi^{\mathrm{far}}_{t}\|_{L^{2}(\Omega)}\approx A_{s}\frac{\|\bm{\rho}\|_{2}}{\sqrt{V}}\sqrt{r_{c}}\,c_{s}^{-1/2}e^{-c_{s}},$ (89)
where
$A_{s}:=\frac{4}{\sqrt{3\pi}}\sqrt{C_{\rho}}\,A_{2}\approx 4.46\sqrt{C_{\rho}}.$ (90)
Assuming that the particle positions are reasonably well distributed (i.e., without significant clustering), we expect $C_{\rho}>1$, but close to $1$, as discussed in Remark 10.
Since $C_{\rho}$ is not known precisely, we treat $A_{s}$ as a mildly adjustable prefactor. Numerical experiments (see Figure 6) indicate that the choice $A_{s}\approx 5$ provides a good fit in practice, corresponding to $\sqrt{C_{\rho}}\approx 1.1$.
@@ 5.2 Aliasing errors due to window function approximations
The sources of error arising from the NUFFT components are analogous to those in fast Ewald summation. In practice, all window functions are truncated. As discussed in Section 4.3.1 for the Gaussian window, the use of the untruncated Fourier transform introduces additional errors in the spreading and interpolation steps. These are commonly referred to as truncation errors. A second source of error arises from aliasing, which is caused by discrete sampling of the window function. Sampling induces a periodization in Fourier space via the Poisson summation formula, leading to overlapping frequency contributions [25].
@@
The coefficients $c_{\bm{k}}(\tilde{\varphi}_{t})$ typically decay in magnitude as $|\bm{k}|$ increases, possibly with oscillations (as for PSWF). However, this decay alone does not guarantee suppression of aliasing errors. Indeed, for the specific choice $\bm{k}=(m/2,0,0)$ and $\bm{r}=(-1,0,0)$, we have $\bm{k}+m\bm{r}=-\bm{k}$. For this choice, since the coefficients are even, we have
$\frac{c^{2}_{\bm{k}+m\bm{r}}(\tilde{\varphi}_{t})}{c^{2}_{\bm{k}}(\tilde{\varphi}_{t})}=1.$ (99)
Thus, in the absence of oversampling, aliasing errors are not uniformly suppressed by the window function alone. Nevertheless, the diagonal scaling in fast Ewald methods—via the $|\hat{f}_{\bm{k}}|^{2}$ factor—provides additional damping.
@@ 5.2.1 A closed-form approximation for the aliasing error
We now derive an approximate model for the aliasing error by selecting the dominant terms in the aliasing sum. This approach is inspired by [19], where a simplified aliasing sum for the relative force error, similar in form to (107) below, is stated.
In (102), the aliasing sum is dominated by the first nonzero terms with $|\bm{r}|=1$, since the PSWF satisfies $|\psi_{0}^{c_{w}}(s)|\leq\psi_{0}^{c_{w}}(1)/s$ for $s\geq 1$ (Lemma 4), implying fast (algebraic) decay of the shifted coefficients as $|\bm{r}|$ increases. These six contributions—corresponding to one shift in each coordinate direction in three dimensions—provide the leading-order term to the aliasing error. We therefore approximate (102) by retaining only these terms:
@@
Since $|\bm{k}|\geq|k_{i}|$, and since $\psi_{0}^{c}$ is even and nonincreasing on $[0,1]$ and bounded by $\psi_{0}^{c}(1)$ on $s>1$ by Lemma 4, we obtain for $\bm{k}=(k_{1},k_{2},k_{3})\in\mathcal{I}_{\bm{m}}\setminus\{\bm{0}\}$,
@@
where the approximation uses that $c_{s}\approx c_{w}$, so that $\psi_{0}^{c_{s}}(2k/m)/\psi_{0}^{c_{w}}(2k/m)\approx 1$ for the relevant range of $k$.
For $k_{i}\in\mathcal{I}_{m}$ and $r_{i}\in\{\pm 1\}$, we have
$\left|\tfrac{2}{m}(k_{i}+mr_{i})\right|\in[1,3]\qquad\text{and}\qquad\left|\tfrac{2}{m}k_{i}\right|\in(0,1].$ (110)
Hence, for $|\bm{r}|=1$, we obtain
@@
Using (109) and (111), and retaining only the axis-aligned modes, we obtain the approximation
$\sum_{\bm{k}\in\mathcal{I}_{\bm{m}}\setminus\{\bm{0}\}}\frac{\psi_{0}^{c_{s}}\!\left(\tfrac{2}{m}|\bm{k}|\right)^{2}\displaystyle\sum_{|\bm{r}|=1}\prod_{i=1}^{3}\psi_{0}^{c_{w}}\!\Big(\tfrac{2}{m}(k_{i}+mr_{i})\Big)^{2}}{|\bm{k}|^{4}\displaystyle\prod_{i=1}^{3}\psi_{0}^{c_{w}}\!\left(\tfrac{2}{m}k_{i}\right)^{2}}$ $\approx 6\,\psi_{0}^{c_{w}}(1)^{2}\sum_{k=1}^{K_{\max}}\frac{1}{k^{4}}\;\lesssim\;\psi_{0}^{c_{w}}(1)^{2}\,\frac{\pi^{4}}{15},$ (112)
@@
Substituting (112) into (107) yields the dimensionless prefactor $C_{\rho}\pi^{2}/15$, which is of order unity and independent of $m$, $L$, and $c_{w}$. We thus arrive at the approximation
$\frac{1}{\sqrt{V}}\|\phi_{t}^{\mathrm{far}}-\phi_{t,h}^{\mathrm{far}}\|_{L^{2}(\Omega)}\approx\frac{\|\bm{\rho}\|_{2}}{\sqrt{V}}\frac{\pi}{\sqrt{15}}\,\sqrt{C_{\rho}L}\,\frac{\psi_{0}^{c_{w}}(1)}{\psi_{0}^{c_{w}}(0)}.$ (113)
Approximating $\psi_{0}^{c_{w}}(1)/\psi_{0}^{c_{w}}(0)\approx A_{2}\,c_{w}^{1/2}e^{-c_{w}}$, with $A_{2}\approx 3.42$ from (145) (Appendix A), we obtain
$\frac{1}{\sqrt{V}}\|\phi_{t}^{\mathrm{far}}-\phi_{t,h}^{\mathrm{far}}\|_{L^{2}(\Omega)}\approx A_{w}\,\frac{\|\bm{\rho}\|_{2}}{\sqrt{V}}\,\sqrt{L}\,c_{w}^{1/2}e^{-c_{w}},$ (114)
where
$A_{w}:=\frac{\pi}{\sqrt{15}}\,\sqrt{C_{\rho}}\,A_{2}\approx 2.8\sqrt{C_{\rho}}.$ (115)
As for the split-model (89), we use the same value of $\sqrt{C_{\rho}}$, treated as a mildly adjustable factor; numerical experiments indicate $\sqrt{C_{\rho}}\approx 1.1$ for finite, well-distributed particle configurations, yielding $A_{w}\approx 3.1$.
@@ 6 Parameter selection
This section presents practical guidelines for selecting the PSWF mollifier and window parameters to achieve a prescribed tolerance $\varepsilon$. The results rely on the error models of Section 5. Unless otherwise noted, we assume neutral, well-distributed particle configurations, no oversampling, and parameters $c_{s},c_{w}$ within the fitted ranges of Table 5.
@@ 6.2 PSWF split parameters
Two parameters control the kernel split: the shape parameter $c_{s}$ and the cutoff $r_{c}$. The cutoff balances the real- and Fourier-space costs, while $c_{s}$ determines the number of Fourier modes. They are related by (see (61))